# LLM+KG’24: The International Workshop on Data Management Opportunities in Unifying Large Language Models + Knowledge Graphs, August 25, 2024, Guangzhou, China

### v1

@@ Abstract
This survey investigates the synergistic relationship between Large Language Models (LLMs) and Knowledge Graphs (KGs), which is crucial for advancing AI’s capabilities in understanding, reasoning, and language processing. It aims to address gaps in current research by exploring areas such as KG Question Answering, ontology generation, KG validation, and the enhancement of KG accuracy and consistency through LLMs. The paper further examines the roles of LLMs in generating descriptive texts and natural language queries for KGs. Through a structured analysis that includes categorizing LLM-KG interactions, examining methodologies, and investigating collaborative uses and potential biases, this study seeks to provide new insights into the combined potential of LLMs and KGs. It highlights the importance of their interaction for improving AI applications and outlines future research directions.
keywords:
[…]
1 Introduction
@@
In the era of artificial intelligence, Large Language Models (LLMs) and Knowledge Graphs (KGs) stand out as two pivotal technologies driving advancements in machine understanding, reasoning, and natural language processing. LLMs, such as OpenAI’s GPT series[1], have demonstrated remarkable capabilities in generating human-like text, answering questions, and even creating content across various domains[2]. On the other hand, KGs organize and integrate information in a structured format, enabling machines to understand and infer relationships between entities in the real world [3]. The synergy between LLMs and KGs opens new frontiers for AI applications, enhancing machines’ ability to process, interpret, and generate information with context and accuracy. Our intention with this survey is to fill the gaps left by previous survey papers. [4, 5, 6, 7, 8], offering fresh perspectives and uncovering new frontiers in the interaction between LLMs and KGs. For instance, our analysis delves deeper into KG Question Answering, encompassing aspects such as multi-hop question answering, multi-hop question generation, and KG-powered chatbots. Additionally, our study extends to KG validation, an area not extensively discussed in previous survey papers. Given the transformative potential and challenges of integrating LLMs with KGs, this paper seeks to explore the following research questions:
1. How can LLMs be utilized to generate descriptive textual information for entities in a KG?
2. How can we employ LLMs in ontology generation?
3. How can LLMs help in detecting inconsistencies within KGs?
@@
4. How can LLMs improve the accuracy, consistency, and completeness of KGs through fact-checking?
@@
5. How can LLMs be effectively employed to generate queries from natural language texts? (Text to Sparql or Cypher)
6. How can LLMs contribute to providing accurate answers for KG Question Answering?
@@
The interplay between LLMs and KGs can be categorized into three distinct types. The first, LLMs for KGs, demonstrates how LLMs are utilized to bolster and refine the functionality of KGs. The second, KG-enhanced LLMs, shows the inverse, where KGs are leveraged to enhance and inform the capabilities of LLMs. Lastly, LLM-KG Cooperation is depicted, representing a collaborative and synergistic relationship where LLMs and KGs work in conjunction to achieve more advanced and complex outcomes. Figure 1 illustrates the categorization of the interaction between LLMs and KGs. The remainder of this paper is structured to include an in-depth exploration of the three distinct types along with their respective sub-categories. The primary focus of this research is to address our formulated research questions. These specific areas of inquiry are highlighted in Figure 1, distinctly marked in pink, mentioned as Tasks Considered in our Research Questions. Those sections highlighted with stars represent topics that were not addressed in previous survey papers. The paper is organized as follows: Section 2 explores the role of LLMs within the context of KGs and Section 3 investigates how KGs can enhance LLMs. In Section 4, we examine the collaborative utilization of LLMs and KGs. Section 5 entails a statistics analysis and the identification of open challenges. Finally, we conclude in Section 6.
Figure 1: The interplay between LLMs and KGs categorization
2 LLM for KG
2.1 KG Construction
KG-to-Text Generation: This segment of the paper is dedicated to addressing Research Question 1, which investigates the effective use of LLMs in generating descriptive textual information for entities within KGs. KG-to-Text generation encompasses the procedure of transforming structured data held within a KG into human-readable, natural-language text [9]. Colas et al. combined a pre-trained LM with graph attention information by modifying the encoder layer. Initially, they convert the KG into a textual string by linearizing its structure. Then, the vector representation becomes contextualized through the incorporation of graph attention data. Several methods involve the fine-tuning of pre-trained language models specifically on KG-to-text datasets to address this particular task [10]. Chen et al. gathered a corpus named KGTEXT which involves matching linked sentences sourced from Wikipedia with knowledge subgraphs obtained from WikiData and then conducting pre-training and subsequently fine-tuning on KG-to-Text datasets. Ke et al. developed a basic yet effective structure-aware semantic aggregation module that integrates into every Transformer layer within the encoder. This module’s purpose is to maintain the structural details present in the input graphs. Then followed by three pre-training tasks aimed at explicitly acquiring graph-text alignments. Li et al. investigated few-shot KG-to-text generation which generates the output with the help of around several hundred instances. Initially, their model aligns KG representations, encoded by graph neural networks, with PLM-based entity representations, aiming to close the semantic gap. It then proposes a specialized strategy, called relation-biased breadth-first search (RBFS), to arrange the KG into a well-structured entity sequence for integration into PLMs. Lastly, the model employs multi-task learning, simultaneously training the main text generation and an auxiliary KG reconstruction task. This approach significantly improves the semantic alignment between the input KG and the resulting text, allowing the model to generate accurate descriptions of the KG.
@@ 2.1 KG Construction
Ontology Creation: Concept Extraction to Validation: This segment of the paper specifically addresses Research Question 2, delving into the application of LLMs in ontology generation for KGs. LLMs exhibit significant promise in the construction of ontologies [14], where the organized representation of knowledge is essential for facilitating intelligent information retrieval and reasoning. LLMs contribute to creating and enriching comprehensive ontologies, with domain-specific fine-tuning enhancing their integration into specialized areas. While constructing a KG, [15] built an ontology for COVID-19 and demonstrated the effectiveness of LLMs. LLMs are promising for ontology generation but face computational intensity and bias challenges, requiring broader evaluation, yet remaining valuable in ontology creation and intelligent systems.
@@
Concept and Relation Extraction: LLMs like GPT-3, with their natural language understanding abilities, are suitable for extracting ontology concepts and relations. Concept extraction is the process of identifying and extracting concepts from text. The process involves extracting linguistic realizations of domain-specific concepts and accumulating semantic term variations with sense disambiguation and domain-specific synonym identification [16]. Relation extraction aims to identify non-taxonomic relations between concepts [17]. It aims to identify relations between text entities and can be enhanced by using techniques like Named Entity Recognition (NER) and additional context provided by LLMs [18, 19].
@@
Property Identification: Research investigates using fine-tuned LLMs for pre-annotating data in ontology property identification, helping annotators efficiently discern relevant properties and their correlation with automated scores, thereby reducing annotation times.[20].
@@
Ontology Enrichment: LLMs can help to improve current ontologies by adding new information or improving concept definitions. This is especially beneficial when working with dynamic domains in which the ontology must adapt to new facts. Advanced machine learning methods, particularly fine-tuned LLMs, have been researched for data pre-annotation and ontology enrichment [20].
@@
Ontology Alignment: LLMs can help to align ontologies by understanding and comparing the semantics of ideas and relationships among ontologies. Ontology alignment with LLMs has been studied in the literature. One method suggests a neurosymbolic design that combines the flexibility of LLMs and the domain orientation of Enterprise KGs (EKGs) [21]. Another work investigates the use of fine-tuned LLMs to pre-annotate data in a lexical extension job, specifically adding descriptive words to an existing ontology [20].
@@
Text-to-Ontology Mapping: Train LLMs to map natural language text to specific ontology concepts and relationships. This can be accomplished using supervised learning, which involves training the model using annotated instances of text-to-ontology mappings. Text-to-ontology mapping using LLMs has been explored in multiple papers. One approach involves using artificial neural networks and classifiers to match texts to relevant ontologies [22].
@@
Ontology Learning: Ontology learning automates the construction of ontologies by identifying concepts from text, aiming for a cost-effective, scalable knowledge representation across disciplines. Recent advances in LLMs, such as BERT, offer a promising alternative to traditional, time-consuming ontology creation, outperforming classic word embedding models in various nlp tasks and providing a more efficient method for knowledge acquisition in ontology development [16, 23, 17, 24].
@@
Entity Extraction and Alignment: The Named Entity Recognition (NER) task involves being provided with a sentence, and the objective is to predict the set of entities present within that sentence[25]. Some papers concentrated on prompting to extract entities. Ashok and Lipton approach named PromptNER, comprises four essential components: a backbone LLM, a prompt outlining the set of entity types, a small set of examples from the target domain, and a specific format for outputting the extracted entities. Some approaches employ instruction-tuning. It is a technique in which pre-trained autoregressive language models undergo finetuning to adhere to natural language instructions and produce responses [26]. Zhou et al. proposed a focused distillation method with instruction tuning tailored to a specific mission.
@@
Relation Extraction: LLMs have notably advanced the complex task of relation extraction between text entities, revolutionizing the identification and interpretation of semantic connections in textual data. In particular, GPT-RE [28] introduced a novel approach by infusing task-specific knowledge and logic-enriched demonstrations. This approach effectively reduces the low relevance issues between entity-relation pairings and overcomes the limitations in explaining input-label correlations. Furthermore, considering the vast array of predefined relation types and the challenges posed by uncontrolled LLMs, the study [29] suggests incorporating a natural language inference module within LLMs. This integration focuses on generating relation triples to greatly improve the usefulness and precision of document-level relation datasets. These advancements in relation extraction demonstrate LLMs’ capability to improve semantic analysis in text, and we have organized these methods by their distinct learning approaches. This classification comprises:
@@
Supervised Fine-tuning known for being prevalent and promising, involves extensively inputting training data to fine-tune LLMs. It allows LLMs to learn and comprehend the structural patterns in data, enhancing their ability to generalize. The Deepstruct model, referenced by [30], uses structure pre-training on diverse corpora to improve language models’ understanding of complex language structures, leading to a more nuanced interpretation of linguistic patterns. Huguet Cabot and Navigli enhanced relation extraction by training a specialized BART-style model [32] with a novel triplet linearization technique, which demonstrates the importance of supervised fine-tuning in advancing LLMs for complex tasks like relation extraction.
@@
Few-shot learning, which relies on a limited number of labeled examples, faces significant challenges in machine learning. The lack of data can cause problems like overfitting and difficulty in understanding complex data relationships, as pointed out by [33]. The advent of LLMs has been a significant breakthrough in overcoming few-shot learning challenges. Xu et al. introduced two strategies for using LLMs in relation extraction: 1) in-context learning (ICL) and 2) data generation. ICL involves using detailed prompts with task definitions and labels to help LLMs understand relation extraction nuances, particularly useful in few-shot settings with limited data. To further address data scarcity, they used LLMs to generate additional training material by creating prompts with instance descriptions and examples. Complementarily, [35] examined the effectiveness of LLMs in few-shot relation extraction through in-context learning. They discovered that LLMs could equal state-of-the-art techniques using just a few examples. Their method included enhancing relation extraction labels with Chain of Thought (CoT) explanations from GPT-3 and fine-tuning the Flan-T5 model with this enriched data.
@@
Zero-shot learning aims to empower models to generalize to new tasks and domains beyond their training, by adapting the pre-training of LLMs to unfamiliar situations. The extensive knowledge embedded in LLMs demonstrates their remarkable capability for zero-shot tasks in various fields. Key studies such as those by [36, 37], and recent advancements [38], exemplify the potential of LLMs in adapting to tasks they were not explicitly trained for. Adding to this discourse, [39] critically assessed ChatGPT’s zero-shot learning capability in temporal relation extraction. The study emphasized ChatGPT’s skill in grasping complex temporal relations but also noted its limitations in consistency and handling long-dependency relations.
2.2 KG Reasoning
KG Reasoning involves deriving new facts from existing ones using logical inference. Recently, there has been a growing emphasis on exploring KG reasoning due to its potential to derive new knowledge and conclusions from existing data[40]. It has applications for KG completion and KG question answering[40]. Luo et al. introduced Reasoning on Graphs (RoG), which combines LLMs and KGs to perform reliable and easily understandable reasoning. To tackle challenges such as hallucinations and knowledge deficiencies, they proposed a framework for planning-retrieval-reasoning. Initially, the planning module produces relation paths in the form of faithful plans using KGs, forming the basis for the RoG approach. Subsequently, these devised plans are employed to extract legitimate paths of reasoning from KGs to carry out reliable reasoning through the retrieval-reasoning module.
2.3 KG Completion
In KG completion, the main tasks are: (i) triple classification, validating the accuracy of a given triple, (ii) link prediction for predicting missing components like the head, tail, or relation in a triple, and (iii) entity classification for categorizing entities, also known as determining the entity’s type[42]. The triple-based methods relying on structural information, are widely used for KG completion, mostly the embedding methods [43, 44, 45]. Several works have attempted to prove more effectiveness than triple-based completion by relying on the extraction of factual knowledge from LMs embedding spaces, instead of simple structures. Examples of these text-based methods, KG-BERT [46], and the fine-tuning of GPT-2 [47]. KG-BERT[46] is a leading example of applying LLMs to KG completion. It fine-tunes for this task by treating triples as textual sequences. The model uses KG-BERT to score sequences of entity-relation-entity, representing these elements through their names or descriptions. These sequences of words are then used as input for fine-tuning the BERT model. KG-BERT was identified with two shortcomings: 1) the model’s struggle to learn relational data in KGs, and 2) difficulty in identifying the correct answers among similar candidates. To improve this, multi-task learning methods combining relation prediction and relevance ranking with link prediction were proposed, as in [48]. This enhanced model better grasps KG relationships and effectively handles lexical similarities. In another study, [49]introduced the Structure-augmented Text Representation (StAR) model, aimed at improving link prediction in KG completion. The model begins with a Siamese-style textual encoder applied to triples, generating two contextualized representations.A distinctive scoring module using dual strategies captures both contextualized and structured knowledge, enhanced by a self-adaptive ensemble scheme with graph embedding techniques to boost performance. Wang et al. presented SimKGC, a straightforward approach to enhance text-based KG completion. It recognizes the crucial aspect of conducting efficient contrastive learning. Building on recent advancements in contrastive learning. SimKGC utilizes a bi-encoder architecture and incorporates three types of negative examples to achieve its objectives. Sequence-to-sequence PLM methods have also shown state-of-the-art performance by redefining the KG completion challenge as a sequence-to-sequence generation task as demonstrated in KG-S2S [51] and GenKGC [52]. GenKGC incorporates innovative techniques such as relation-guided demonstration and entity-aware hierarchical decoding, aimed at improving the quality of representation learning and accelerating the inference process. Most of these recent text-based methods require specific fine-tuning for KGs, which limits their efficiency. To address this limitation, new training free methods with no fine-tuning were recently proposed, as in KICGPT [53].
2.4 KG Embedding
KG Embedding involves mapping elements of a KG into vectors in a continuous vector space. Unlike Graph Neural Networks, which learn KG embeddings from data explicitly represented by nodes and edges, LM learns KG embeddings from natural language. Given a piece of text, LMs learn representations of entities, which capture context information containing relations of entities. The entity relations are found by an attention mechanism, which can create an association between two entities independent of their location. The more text an LM sees the more entities and their relationships the LM can learn; the larger an LM the more contextual information the representation contains. LLMs contain billions of parameters and are trained over a huge amount of textual data (e.g. 45 TB for GPT-3). LLMs perfectly meet these requirements and so are a powerful tool for learning KG embeddings. LLMs’ ability to KG embedding attracts researchers to use them directly to conduct KG tasks, like entity alignment [54]. We can also use the representation of entities learned by LLMs in the small-sized models and this should significantly reduce the amount of training data needed and the time of training and meanwhile create well-generalized models. An extensive experiment is needed to investigate the efficiency of applying embeddings of LLMs into small-sized models for KG analysis tasks.
2.5 KG Validation
2.5.1 Fact Checking
In this section, we address Research Question 4, which is centered on how LLMs can enhance the accuracy, consistency, and completeness of KGs through the process of fact-checking. The advent of LLMs has the potential to significantly impact the fight against misinformation, presenting both opportunities and challenges. LLMs, with their extensive world knowledge and strong reasoning abilities, offer promise in combating misinformation. However, the flip side is that these models can also be easily manipulated to generate deceptive misinformation at scale [55]. An approach to address misinformation in a KG involves verbalizing each triple within the KG and prompting LLMs to identify instances of misinformation [56, 1]. To address the limitations of potentially outdated or insufficient knowledge in LLMs for detecting factual errors, some research has explored augmenting LLMs with external knowledge[57] or utilizing tools [58] for misinformation detection. The methodologies developed for factual checking in LLMs can be applied to KG fact-checking by converting KG triples into textual representations.
2.5.2 Inconsistency Detection

### v2

# Research Trends for the Interplay between Large Language Models and Knowledge Graphs

@@ Abstract.
This survey investigates the synergistic relationship between Large Language Models (LLMs) and Knowledge Graphs (KGs), which is crucial for advancing AI’s understanding, reasoning, and language processing capabilities. It aims to address gaps in current research by exploring areas such as KG Question Answering, ontology generation, KG validation, and enhancing KG accuracy and consistency through LLMs. The paper further examines the roles of LLMs in generating descriptive texts and natural language queries for KGs. Through a structured analysis that includes categorizing LLM-KG interactions, examining methodologies, and investigating collaborative uses and potential biases, this study seeks to provide new insights into the combined potential of LLMs and KGs. It highlights the importance of their interaction for improving AI applications and outlines future research directions.
VLDB Workshop Reference Format:
VLDB 2024 Workshop: LLM+KG.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment. ISSN 2150-8097.
1. Introduction
@@
In the era of artificial intelligence, Large Language Models (LLMs) and Knowledge Graphs (KGs) stand out as two pivotal technologies driving advancements in machine understanding, reasoning, and natural language processing. LLMs, such as OpenAI’s GPT series (Bang et al., 2023), have demonstrated remarkable capabilities in generating human-like text, answering questions, and even creating content across various domains (Khorashadizadeh et al., 2023). On the other hand, KGs organize and integrate information in a structured format, enabling machines to understand and infer relationships between entities in the real world (Xu et al., 2021). The synergy between LLMs and KGs opens new frontiers for AI applications, enhancing machines’ ability to process, interpret, and generate information with context and accuracy. Our intention with this survey is to fill the gaps left by previous survey papers (Pan et al., 2024, 2023; Hu et al., 2023; Yang et al., 2024), offering fresh perspectives and uncovering new frontiers in the interaction between LLMs and KGs. For instance, our analysis delves deeper into KG Question Answering, encompassing multi-hop question answering, multi-hop question generation, and KG-powered chatbots. Additionally, our study extends to KG validation, an area not extensively discussed in previous survey papers. Table 1 outlines the categories discussed in previous research papers and this survey paper. Given the transformative potential and challenges of integrating LLMs with KGs, this paper seeks to explore the following research questions:
  1. 1.
    How can LLMs generate descriptive textual information for entities in a KG?
  2. 2.
    How can we employ LLMs in ontology generation?
  3. 3.
    How can LLMs help in detecting inconsistencies within KGs?
  4. 4.
@@
    How can LLMs improve the accuracy, consistency, and completeness of KGs through fact-checking?
  5. 5.
    How can LLMs contribute to providing accurate answers for KG Question Answering?
  6. 6.
@@
    How can LLMs effectively generate queries from natural language texts? (Text to Sparql or Cypher)
@@
The interplay between LLMs and KGs can be categorized into three distinct types. The first, LLMs for KGs, demonstrates how LLMs are utilized to bolster and refine the functionality of KGs. The second, KG-enhanced LLMs, shows the inverse, where KGs are leveraged to enhance and inform the capabilities of LLMs. Lastly, LLM-KG Cooperation is depicted, representing a collaborative and synergistic relationship where LLMs and KGs work together to achieve more advanced and complex outcomes. Figure 1 illustrates the categorization of the interaction between LLMs and KGs. The remainder of this paper is structured to include an in-depth exploration of the three distinct types along with their respective sub-categories. The primary focus of this research is to address our formulated research questions. These specific areas of inquiry are highlighted in Figure 1, distinctly marked in pink, mentioned as Tasks Considered in our Research Questions. Those sections highlighted with stars represent topics that were not addressed in previous survey papers. The paper is organized as follows: Section 2 explores the role of LLMs within the context of KGs and Section 3 investigates how KGs can enhance LLMs. In Section 4, we examine the collaborative utilization of LLMs and KGs. Section 5 entails a statistics analysis and the identification of open challenges. Finally, we conclude in Section 6.
Figure 1. Categorization of the interplay between LLMs and KGs
Table 1. Categorizations addressed by previous survey papers
Main Category Subcategory Survey: (Pan et al., 2024) Survey: (Pan et al., 2023) Survey: (Hu et al., 2023) Survey: (Yang et al., 2024) Our Survey KG Construction Relation and Attribute Extraction Entity Extraction and Alignment Event Detection or Extraction Ontology Creation KG-to-Text Generation KG-to-Text Generation KG Reasoning KG Reasoning KG Completion Entity, Relation and Triple Classification Entity Prediction Relation Prediction KG Embedding KG Embedding KG-enhanced LLM KG-enhanced LLM KG Validation Fact Checking Inconsistency Detection KG Question Answering Complex Question Answering Multi-Hop Question Generation Knowledge Graph Chatbots Query Generation from natural text Querying Large Language Models with SPARQL
2. LLM for KG
2.1. KG Construction
2.1.1. Ontology Creation
@@
This section of the paper explicitly addresses Research Question 2, delving into the application of LLMs in ontology generation for KGs. LLMs exhibit significant promise in constructing ontologies (Neuhaus, 2023), where the organized representation of knowledge is essential for facilitating intelligent information retrieval and reasoning. LLMs contribute to creating and enriching comprehensive ontologies, with domain-specific fine-tuning enhancing their integration into specialized areas. While constructing a KG, (Ezzabady et al., 2024) built an ontology for COVID-19 and demonstrated the effectiveness of LLMs. LLMs are promising for ontology generation but face computational intensity and bias challenges. They require broader evaluation yet remain valuable in ontology creation and intelligent systems.
Concept and Relation Extraction:
@@
LLMs like GPT-3, with their natural language understanding abilities, are suitable for extracting ontology concepts and relations. Concept extraction is the process of identifying and extracting concepts from text. The process involves extracting linguistic realizations of domain-specific concepts and accumulating semantic term variations with sense disambiguation and domain-specific synonym identification (Schaeffer et al., 2023). Relation extraction aims to identify non-taxonomic relations between concepts (Babaei Giglou et al., 2023). It seeks to identify relations between text entities and can be enhanced by using techniques like Named Entity Recognition (NER) and additional context provided by LLMs (Caufield et al., 2023; Chang, 2023).
Property Identification:
@@
Contributions like (Straková et al., 2023) investigate using fine-tuned LLMs for pre-annotating data in ontology property identification, helping annotators efficiently discern relevant properties and their correlation with automated scores, thereby reducing annotation times.
Ontology Enrichment:
@@
LLMs can help to improve current ontologies by adding new information or improving concept definitions. This is especially beneficial when working with dynamic domains in which the ontology must adapt to new facts. Advanced machine learning methods, remarkably fine-tuned LLMs, have been researched for data pre-annotation and ontology enrichment (Straková et al., 2023).
Ontology Alignment:
@@
LLMs can help to align ontologies by understanding and comparing the semantics of ideas and relationships among ontologies. Ontology alignment with LLMs has been studied in the literature. One method suggests a neurosymbolic design that combines the flexibility of LLMs and the domain orientation of Enterprise KGs (EKGs) (Baldazzi et al., 2023). Another work investigates the use of fine-tuned LLMs to pre-annotate data in a lexical extension job, specifically adding descriptive words to an existing ontology (Straková et al., 2023).
Text-to-Ontology Mapping:
@@
Train LLMs to map natural language text to specific ontology concepts and relationships. This can be accomplished using supervised learning, which involves training the model using annotated instances of text-to-ontology mappings. Text-to-ontology mapping using LLMs has been explored in multiple papers. One approach involves using artificial neural networks and classifiers to match texts to relevant ontologies (Korel et al., 2023).
Ontology Learning:
@@
Ontology learning automates the construction of ontologies by identifying concepts from text, aiming for a cost-effective, scalable knowledge representation across disciplines. Recent advances in LLMs, such as BERT, offer a promising alternative to traditional, time-consuming ontology creation, outperforming classic word embedding models in various nlp tasks and providing a more efficient method for knowledge acquisition in ontology development (Schaeffer et al., 2023; Chen et al., 2023; Babaei Giglou et al., 2023; Funk et al., 2023).
2.1.2. Entity Extraction and Alignment:
@@
The Named Entity Recognition (NER) task involves being provided with a sentence, and the objective is to predict the set of entities present within that sentence(Ashok and Lipton, 2023). Some papers concentrated on prompting to extract entities. Ashok and Lipton approach named PromptNER, comprises four essential components: a backbone LLM, a prompt outlining the set of entity types, a small set of examples from the target domain, and a specific format for outputting the extracted entities. Some approaches employ instruction-tuning. It is a technique in which pre-trained autoregressive language models undergo finetuning to adhere to natural language instructions and produce responses (Ouyang et al., 2022). Zhou et al. proposed a focused distillation method with instruction tuning tailored to a specific mission.
2.1.3. Relation Extraction:
@@
LLMs have notably advanced the complex task of relation extraction between text entities, revolutionizing the identification and interpretation of semantic connections in textual data. In particular, GPT-RE (Wan et al., 2023) introduced a novel approach by infusing task-specific knowledge and logic-enriched demonstrations. This approach effectively reduces the low relevance issues between entity-relation pairings and overcomes the limitations in explaining input-label correlations. Furthermore, considering the vast array of predefined relation types and the challenges posed by uncontrolled LLMs, the study (Li et al., 2023b) suggests incorporating a natural language inference module within LLMs. This integration focuses on generating relation triples to greatly improve the usefulness and precision of document-level relation datasets. These advancements in relation extraction demonstrate LLMs’ capability to improve semantic analysis in text, and we have organized these methods by their distinct learning approaches. This classification comprises:
Supervised Fine-tuning:
@@
Supervised fine-tuning known for being prevalent and promising, involves extensively inputting training data to fine-tune LLMs. It allows LLMs to learn and comprehend the structural patterns in data, enhancing their ability to generalize. The Deepstruct model, referenced by (Wang et al., 2023a), uses structure pre-training on diverse corpora to improve language models’ understanding of complex language structures, leading to a more nuanced interpretation of linguistic patterns. Huguet Cabot and Navigli enhanced relation extraction by training a specialized BART-style model (Lewis et al., 2020) with a novel triplet linearization technique, which demonstrates the importance of supervised fine-tuning in advancing LLMs for complex tasks like relation extraction.
Few-shot learning:
@@
Few-shot learning, which relies on a limited number of labeled examples, faces significant challenges in machine learning. The lack of data can cause problems like overfitting and difficulty in understanding complex data relationships, as pointed out by (Huang et al., 2020). The advent of LLMs has been a significant breakthrough in overcoming few-shot learning challenges. Xu et al. introduced two strategies for using LLMs in relation extraction: 1) in-context learning (ICL) and 2) data generation. ICL involves using detailed prompts with task definitions and labels to help LLMs understand relation extraction nuances, which is particularly useful in few-shot settings with limited data. To further address data scarcity, they used LLMs to generate additional training material by creating prompts with instance descriptions and examples. In a complementary study, (Wadhwa et al., 2023) examined the effectiveness of LLMs in few-shot relation extraction through in-context learning. They discovered that LLMs could equal state-of-the-art techniques using just a few examples. Their method included enhancing relation extraction labels with Chain of Thought (CoT) explanations from GPT-3 and fine-tuning the Flan-T5 model with this enriched data.
Zero-shot learning:
@@
Zero-shot learning aims to empower models to generalize to new tasks and domains beyond their training by adapting the pre-training of LLMs to unfamiliar situations. The extensive knowledge embedded in LLMs demonstrates their remarkable capability for zero-shot tasks in various fields. Key studies such as those by (Kojima et al., 2023; Wei et al., 2023a), and recent advancements (Li et al., 2023c), exemplify the potential of LLMs in adapting to tasks they were not explicitly trained for. Adding to this discourse, (Yuan et al., 2023) critically assessed ChatGPT’s zero-shot learning capability in temporal relation extraction. The study emphasized ChatGPT’s skill in grasping complex temporal relations but also noted its limitations in consistency and handling long-dependency relations.
2.2. KG-to-Text Generation: